OpenStudy (anonymous):

Simplify the expression 3 + 9 × 4 ÷ 2.S elect one of the options below as your answer: A. 24 B. 21 C. 16 D. 19

OpenStudy (angela210793):

9*4=36 36/2=18 18+3=21 So B

jhonyy9 (jhonyy9):

3/9=1/3 x4=4/3 /10/4 =4/3 x 4/10 =16/30 =8/15

OpenStudy (angela210793):

noo Jhonny u first multiply then divide then add

jhonyy9 (jhonyy9):

sorry but i see that 3 is divided with 9

OpenStudy (angela210793):

no that's a plus

jhonyy9 (jhonyy9):

ty